# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v1

@@ 4.2 Fusion Mechanisms
Attention-based fusion dynamically weights cross-modal interactions for task-specific reasoning. RAMM Yuan et al. (2023) employs a dual-stream co-attention transformer, combining self-attention (intra-modal) and cross-attention (inter-modal) to fuse retrieved biomedical images/texts with input data. Similarly, EMERGE Zhu et al. (2024b) integrates Electronic Health Records (EHR) data with a cross-attention adaptive network, while MORE Cui et al. (2024) uses cross-attention to weigh multimodal results for commonsense reasoning. For user-centric tasks, RAGTrans Cheng et al. (2024) integrates user-aware attention to fuse social media features, and AlzheimerRAG Lahiri and Hu (2024) applies cross-modal attention to synthesize text, images, and tables. Video-text alignment is advanced by MV-Adapter Jin et al. (2024), which leverages a Cross Modality Tying (CMT) to align embeddings, and M2-RAAP Dong et al. (2024b), enhances video-text fusion using an Auxiliary-Caption-Guided (ACG) strategy. It employs Frame-Caption Re-weighting (FCR) to boost frames with high inter-modal consistency and Text-Caption Re-weighting (TCR) to adjust weights based on intra-modal similarity. A Mutual-guided (Mug) Alignment Head then filters misaligned features via a dot-product similarity and frame-to-token attention, producing refined frame-specific text representations. Generative refinement is demonstrated in Xu et al. (2024a), where gated cross-attention conditions text generation on visual features. Similarly, C3Net Zhang et al. (2024c) enhances alignment between queries and candidates within CLIP’s contrastive space. Mu-RAG Chen et al. (2022a) uses intermediate fusion with cross-attention for open QA, balancing modality-specific and joint representations. Kim et al. (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
Unified Frameworks and Projections
@@
Architectures unifying multimodal inputs through frameworks or projections address scalability. Su et al. (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al. (2024b) iteratively consolidates multimodal results into knowledge representations. For document understanding, M3DocRAG Cho et al. (2024) flattens multi-page documents into a unified embedding tensor, and PDF-MVQA Ding et al. (2024c) combines Region-of-Interest (RoI)-based (VisualBERT) and patch-based (CLIP) VLPMs Long et al. (2022). DQU-CIR Wen et al. (2024) exemplifies raw-data unification by converting images into text captions for complex queries and overlaying text onto images for simple queries. The model then fuses embeddings using MLP-learned weights. SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text representation for subsequent processing. UFineBench Zuo et al. (2024) uses a shared granularity decoder for ultra-fine text-person retrieval. Nguyen et al. (2024) introduce Dense2Sparse (D2S) projection, converting dense embeddings from models like BLIP/ALBEF Li et al. (2022a) into sparse lexical vectors via layer normalization and probabilistic expansion control, optimizing storage and interpretability.
@@ 4.3 Augmentation Techniques
Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs. To address this, augmentation techniques have been introduced to enhance the utility of retrieved data before generation, ensuring that the model can better interpret, structure, and integrate multimodal information Gao et al. (2023).
Context Enrichment
@@
This focuses on enhancing the relevance of retrieved knowledge for the generation phase by refining external knowledge or retrieved data. General approaches such as EMERGE Zhu et al. (2024b) integrate entity relationships and semantic descriptions to provide richer context. MiRAG Adjali et al. (2024) improves retrieval effectiveness by expanding initial queries through entity retrieval and query reformulation, enhancing subsequent stages for the visual question-answering task. Wiki-LLaVA Caffagni et al. (2024) incorporates text chunks and image tokens into multimodal inputs to enrich prompts for generation. Video-RAG Luo et al. (2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract additional auxiliary multimodal context. Img2Loc Zhou et al. (2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions. Finally, Xue et al. (2024) introduce semantic-enhanced prompts, aligning structured image data with queries for contextual accuracy.
Adaptive and Iterative Retrieval
For more complex queries, dynamic retrieval mechanisms have proven effective. SKURG Yang et al. (2023) employs adaptive retrieval steps to dynamically adjust the number of hops based on the query complexity. IRAMIG Liu et al. (2024b) iteratively refines multimodal queries by incorporating feedback from previous iterations and context created by the fusion of retrieved items from text, images, and audio modalities. Similarly, OMG-QA Nan et al. (2024) introduces a multi-round retrieval strategy, where each retrieval step incorporates episodic memory to refine subsequent queries. An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback. Taking a different approach, SAM-RAG Zhai (2024) further improves relevance by dynamically selecting the optimal number of retrieved contexts for each query. MMed-RAG Xia et al. (2024a) uses context filtering to discard low-relevance results. OmniSearch Li et al. (2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content. mR2AG Zhang et al. (2024g) introduces reflection mechanisms to enhance retrieval effectiveness. Retrieval-Reflection dynamically determines when external knowledge is necessary, while Relevance-Reflection filters retrieved content, retaining only task-critical evidence. Similarly, RAGAR Khaliq et al. (2024) iteratively refines retrieved data by leveraging prior responses and multimodal analysis, ensuring contextual consistency.
@@ 4.4 Generation Techniques
In-Context Learning
@@
In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining. Models such as RMR Tan et al. (2024), Sharifymoghaddam et al. (2024), and RA-CM3 Yasunaga et al. (2023), extend this paradigm to multimodal RAG settings. RAG-Driver Yuan et al. (2024) refines ICL by retrieving relevant driving experiences from a memory database, ensuring scenario-specific contextual alignment. MSIER Luo et al. (2024a) further enhances example selection through a Multimodal Supervised In-Context Examples Retrieval framework, leveraging a foundation MLLM scorer to evaluate both textual and visual relevance. Meanwhile, Raven Rao et al. (2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
Reasoning
@@
Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems. RAGAR Khaliq et al. (2024) introduces Chain of RAG and Tree of RAG to iteratively refine fact-checking queries and explore branching reasoning paths for more robust evidence generation. VisDoM Suri et al. (2024) integrates CoT with evidence curation to ensure logical and contextual accuracy, while SAM-RAG Zhai (2024) employs reasoning chains alongside multi-stage answer verification to enhance the relevance, utility, and support of generated responses. Meanwhile, LDRE Yang et al. (2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
Instruction Tuning
Several works have fine-tuned generation components for specific applications. RA-BLIP Ding et al. (2024b), uses Q-Former architecture from InstructBLIP Dai et al. (2023) to extract visual features based on question instructions. RAGPT Lang et al. (2025) uses a context-aware prompter, which captures contextual knowledge from relevant instances and generates dynamic prompts. mR2AG Zhang et al. (2024g) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks. RagVL Chen et al. (2024d) employs instruction tuning to enhance the ranking capability of MLLMs, serving them as a re-ranker for filtering the top-k retrieved images. Jang et al. (2024) propose a Visual Delta Generator to discriminate differences between image pairs and produce a textual response. MMed-RAG Xia et al. (2024a) uses preference fine-tuning to help the model learn when to rely on retrieved information and when to trust its internal knowledge. Additionally, several methods have been proposed to improve generation quality by constructing datasets from the previous mistakes of LLMs and fine-tuning them accordingly. MegaPairs Zhou et al. (2024a) synthesizes large-scale multimodal instruction-tuning datasets for this purpose. Similarly, Surf Sun et al. (2024a) creates a dataset of positive and negative pairs derived from the prior errors of LVLMs and subsequently fine-tunes the LVLM using this dataset. Additionally, Rule Xia et al. (2024b) refines Med-LVLM by applying direct preference optimization on curated preference data, specifically focusing on instances where overreliance on retrieved contexts resulted in errors.
Source Attribution and Evidence Transparency
@@
Ensuring source attribution in multimodal RAG systems is a key focus of recent research. MuRAR Zhu et al. (2024d) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring coherence and informativeness. VISA Ma et al. (2024b) uses large vision-language models (VLMs) to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots. Similarly, OMG-QA Nan et al. (2024) enhances transparency by prompting the LLM to explicitly cite evidence in generated responses.
@@ 4.5 Training Strategies
Training multimodal RAG models involves a multi-stage process to effectively handle cross-modal interactions Chen et al. (2022a). Pretraining establishes the foundation using large paired datasets to learn cross-modal relationships while fine-tuning adapts models to specific tasks by leveraging cross-modal attention Ye et al. (2019). For instance, REVEAL Hu et al. (2023) integrates multiple training objectives. Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where the model predicts text continuations from a prefix and an image. Supporting losses include Contrastive Loss ($L_{\text{contra}}$) for aligning queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to improve embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to align query and knowledge embeddings. During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning.
Alignment
@@
The eCLIP loss Kumar and Marttinen (2024) extends contrastive training by integrating expert-annotated data and an auxiliary Mean Squared Error (MSE) loss to refine embedding quality. Mixup strategies further augment training by generating synthetic positive pairs, improving generalization in contrastive learning Kumar and Marttinen (2024). Dense2Sparse Nguyen et al. (2024) incorporates two unidirectional losses: an image-to-caption loss $\ell(I\to C)$ and a caption-to-image loss $\ell(C\to I)$. It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations. Details of formulas for widely used RAG loss functions can be found in Appendix D.
Robustness
@@
Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024). Managing noisy retrieval inputs is critical for maintaining model performance. MORE Cui et al. (2024) injects irrelevant results during training to enhance focus on relevant inputs. AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment. RAGTrans Cheng et al. (2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information. RA-BLIP Ding et al. (2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning. This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters. RagVL Chen et al. (2024d) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise. Finally, RA-CM3 Yasunaga et al. (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, ver/.style= fill=gray!10, rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=8em , leaf/.style= text opacity=1, inner sep=2pt, fill opacity=.5, fill=hidden-blue!90, text=black, text width=44.5em font=, inner xsep=0.4em, inner ysep=0.6em, draw, , , [Multimodal RAG, ver [ Retrieval
Strategies (§4.1) [ Efficient-Search and
Similarity Retrieval
 (§4.1) [ Maximum Inner Product Search (MIPS) [ TPU-KNN Chern et al. (2022), ScaNN Guo et al. (2020), MAXSIM score Chan and Ng (2008), OmniSearch Li et al. (2024b), MuRAG Chen et al. (2022a), RetrievalAttention Liu et al. (2024a), RA-CM3 Yasunaga et al. (2023), RAVEN Rao et al. (2024) FactMM-RAG Sun et al. (2024b) , leaf, text width=44.5em ] ] [ Multimodal
Encoders [ CLIP Radford et al. (2021), BLIP Li et al. (2022a), MARVEL Zhou et al. (2024c), ALIGN Jia et al. (2021), FLAVA Singh et al. (2022), UniVL-DR Liu et al. (2023b), UniIR Wei et al. (2025), GME Zhang et al. (2024j), VISTA Zhou et al. (2024b), ColPali Faysse et al. (2024), InternVideo Wang et al. (2022), Ovis Lu et al. (2024) Mi-RAG Adjali et al. (2024) , leaf, text width=44.5em ] ] ] [ Modality-Centric
Retrieval (§4.1) [ Text-Centric [ Contriever Izacard et al. (2022), GTE Li et al. (2023c), Re-Imagen Chen et al. (2022b), BM25 Robertson and Zaragoza (2009), BGE-1.5 Xiao et al. (2023), BGE-M3 Chen et al. (2024a), CapRet Shohan et al. (2024), OMG-QA Nan et al. (2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al. (2024b), RAFT Zhang et al. (2024i), CRAG Yan et al. (2024), M2-RAG Ma et al. (2024c) , leaf, text width=44.5em ] ] [ Vision-Centric [ VQA4CIR Feng et al. (2023), Unifashion Zhao et al. (2024), Jang et al. Jang et al. (2024), Pic2word Saito et al. (2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al. (2023), Joshi et al. Joshi et al. (2024), VISA Ma et al. (2024b), ImgRet Shohan et al. (2024), EchoSight Yan and Xie (2024), Xue et al. Xue et al. (2024) , leaf, text width=44.5em ] ] [ Video-Centric [ iRAG Arefeen et al. (2024), VideoRAG Ren et al. (2025), VideoRAG Jeong et al. (2025), MV-Adapter Jin et al. (2024), OmAgent Zhang et al. (2024f), CM2 Kim et al. (2024), Video-RAG Luo et al. (2024b), CTCH Shen et al. (2024), RTime Du et al. (2024), VideoMAE Tong et al. (2022), DrVideo Ma et al. (2024d) , leaf, text width=44.5em ] ] [ Document Retrieval [ ColPali Faysse et al. (2024), ColQwen2 Wang et al. (2024b), M3DocVQA Cho et al. (2024), ViTLP Mao et al. (2024), DocLLM Wang et al. (2024a), CREAM Zhang et al. (2024b), mPLUG-DocOwl 1.5 Hu et al. (2024a), mPLUG-DocOwl 2 Hu et al. (2024b), VisDom Suri et al. (2024), DSE Ma et al. (2024a) , leaf, text width=44.5em ] ] ] [ Re-ranking
Strategies (§4.1) [ Optimized Example
Selection [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024b), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=44.5em ] ] [ Relevance Score
Evaluation [ RAG-Check Mortaheb et al. (2025a, b), UniRaG Zhi Lim et al. (2024), MR2AG Zhang et al. (2024h), LDRE Yang et al. (2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al. (2024), OMG-QA Nan et al. (2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al. (2024a) , leaf, text width=44.5em ] ] [ Filtering
Mechanisms [ MAIN-RAG Chang et al. (2024), MM-Embed Lin et al. (2024a), GME Zhang et al. (2024j), Img2Loc Zhou (2024), MuRAR Zhu et al. (2024d) RAFT Zhang et al. (2024i) , leaf, text width=44.5em ] ] ] ] [ Fusion
Mechanisms (§4.2) [ Score Fusion
and Alignment
 (§4.2) [ M3 Cai et al. (2025)  Zhi Lim et al. (2024),  Sharifymoghaddam et al. (2024), REVEAL Hu et al. (2023), RAG-Driver Yuan et al. (2024), LLM-RA Jian et al. (2024),  Riedler and Langer (2024), VISA Ma et al. (2024b), MA-LMM He et al. (2024),  Xue et al. (2024), RA-BLIP Ding et al. (2024b), Re-IMAGEN Chen et al. (2022b), MegaPairs Zhou et al. (2024a), Wiki-LLaVA Caffagni et al. (2024), VISRAG Yu et al. (2024) , leaf, text width=55.0em ] ] [ Attention-Based
Mechanisms (§4.2) [ RAMM Yuan et al. (2023), EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), RAGTrans Cheng et al. (2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al. (2024), C3Net Zhang et al. (2024c),  Xu et al. (2024a),  Kim et al. (2024), M2-RAAP Dong et al. (2024b), Mu-RAG Chen et al. (2022a), Ou et al. Ou et al. (2025), CADMR Khalafaoui et al. (2024) , leaf, text width=55.0em ] ] [ Unified Frameworks
and Projections
 (§4.2) [ Hybrid-RAG Su et al. (2024b), Dense2Sparse Nguyen et al. (2024), IRAMIG Liu et al. (2024b), M3DocRAG Cho et al. (2024), DQU-CIR Wen et al. (2024), PDF-MVQA Ding et al. (2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al. (2024),  Li et al. (2022a), , leaf, text width=55.0em ] ] ] [ Augmentation
Techniques (§4.3) [ Context-Enrichment (§4.3) [ EMERGE Zhu et al. (2024b), MiRAG Adjali et al. (2024), Wiki-LLaVA Caffagni et al. (2024), Video-RAG Luo et al. (2024b), Img2Loc Zhou et al. (2024e),  Xue et al. (2024) , leaf, text width=55.0em ] ] [ Adaptive and
Iterative
Retrieval (§4.3) [ SKURG Yang et al. (2023), IRAMIG Liu et al. (2024b), OMG-QA Nan et al. (2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al. (2024a), OmniSearch Li et al. (2024b), mR2AG Zhang et al. (2024g), RAGAR Khaliq et al. (2024) , leaf, text width=55.0em ] ] ] [ Generation
Techniques (§4.4) [ In-Context
Learning (§4.4) [ RMR Tan et al. (2024),  Sharifymoghaddam et al. (2024), RA-CM3 Yasunaga et al. (2023), RAG-Driver Yuan et al. (2024), MSIER Luo et al. (2024a), Raven Rao et al. (2024) , leaf, text width=55.0em ] ] [ Reasoning (§4.4) [ RAGAR Khaliq et al. (2024), VisDoM Suri et al. (2024), SAM-RAG Zhai (2024), LDRE Yang et al. (2024) , leaf, text width=55.0em ] ] [ Instruction
Tuning (§4.4) [ RA-BLIP Ding et al. (2024b), RAGPT Lang et al. (2025), mR2AG Zhang et al. (2024g), RagVL Chen et al. (2024d), Jang et al. (2024), MMed-RAG Xia et al. (2024a), MegaPairs Zhou et al. (2024a), Surf Sun et al. (2024a), Rule Xia et al. (2024b) , leaf, text width=55.0em ] ] [ Source
Attribution (§4.4) [ MuRAR Zhu et al. (2024d), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=55.0em ] ] ] [ Training Strategies (§4.5) [ Alignment (§4.5) [ VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al. (2024),  Zhi Lim et al. (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al. (2024) , leaf, text width=55.0em ] ] [ Robustness (§4.5) [  Buettner and Kovashka (2024), MORE Cui et al. (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al. (2024), RA-BLIP Ding et al. (2024b), RagVL Chen et al. (2024d), RA-CM3 Yasunaga et al. (2023) , leaf, text width=55.0em ] ] ] ] ]
Figure 2: Taxonomy of recent advances and enhancements in multimodal retrieval-augmented generation research.

### v2

@@ 4.2.2 Attention-Based Mechanisms
For video-text alignment, MV-Adapter Jin et al. (2024) leverages Cross Modality Tying to align embeddings, and M2-RAAP Dong et al. (2024b) enhances fusion through an auxiliary caption-guided strategy that re-weights frames and text captions based on intra-modal similarity. A mutual-guided alignment head then filters misaligned features using dot-product similarity and frame-to-token attention, generating refined frame-specific text representations. Xu et al. (2024a) condition text generation on visual features using gated cross-attention, and Mu-RAG Chen et al. (2022a) employs intermediate cross-attention for open-domain QA. Kim et al. (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
4.2.3 Unified Frameworks and Projections
Unified frameworks and projection methods consolidate multimodal inputs into coherent representations. Su et al. (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al. (2024b) iteratively integrates multimodal results into unified knowledge representations. M3DocRAG Cho et al. (2024) flattens multi-page documents into a single embedding tensor, and PDF-MVQA Ding et al. (2024c) fuses Region-of-Interest (RoI)-based and patch-based (CLIP) vision-language models Long et al. (2022).
@@ 4.2.3 Unified Frameworks and Projections
DQU-CIR Wen et al. (2024) unifies raw data by converting images into text captions for complex queries and overlaying text onto images for simple ones, then fusing embeddings via MLP-learned weights. SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text for subsequent processing. UFineBench Zuo et al. (2024) utilizes a shared granularity decoder for ultra-fine text–person retrieval. Nguyen et al. (2024) introduce Dense2Sparse projection, converting dense embeddings from models like BLIP/ALBEF Li et al. (2022a) into sparse lexical vectors using layer normalization and probabilistic expansion control to optimize storage and interpretability.
@@ 4.3 Augmentation Techniques
Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs. To address this, augmentation techniques have been introduced to refine retrieved data beforehand, improving multimodal interpretation, structuring, and integration Gao et al. (2023).
4.3.1 Context Enrichment
@@
This focuses on enhancing the relevance of retrieved knowledge by refining or expanding retrieved data. General approaches incorporate additional contextual elements (e.g., text chunks, image tokens, structured data) to provide a richer grounding for generation Caffagni et al. (2024); Xue et al. (2024). EMERGE Zhu et al. (2024b) enriches context by integrating entity relationships and semantic descriptions. MiRAG Adjali et al. (2024) expands initial queries through entity retrieval and reformulation, enhancing subsequent stages for the visual question-answering. Video-RAG Luo et al. (2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract auxiliary multimodal context. Img2Loc Zhou et al. (2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions.
4.3.2 Adaptive and Iterative Retrieval
For more complex queries, dynamic retrieval mechanisms have proven effective. Adaptive retrieval approaches optimize relevance by adjusting retrieval dynamically. SKURG Yang et al. (2023) determines the number of retrieval hops based on query complexity. SAM-RAG Zhai (2024) and mR2AG Zhang et al. (2024f) dynamically assess the need for external knowledge and filter irrelevant content using MLLMs to retain only task-critical information. MMed-RAG Xia et al. (2024a) further improves retrieval precision by discarding low-relevance results, while OmniSearch Li et al. (2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content.
Iterative approaches refine results over multiple steps by incorporating feedback from prior iterations. IRAMIG Liu et al. (2024b) improves multimodal retrieval by dynamically updating queries based on retrieved content. Similarly, OMG-QA Nan et al. (2024) introduces a multi-round retrieval strategy, where each retrieval step incorporates episodic memory to refine subsequent queries. An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback. RAGAR Khaliq et al. (2024) further enhances contextual consistency by iteratively adjusting retrieval based on prior responses and multimodal analysis.
@@ 4.4 Generation Techniques
4.4.1 In-Context Learning
In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining. Models such as RMR Tan et al. (2024), Sharifymoghaddam et al. (2024), and RA-CM3 Yasunaga et al. (2023), extend this paradigm to multimodal RAG settings.
@@ 4.4.1 In-Context Learning
RAG-Driver Yuan et al. (2024) refines ICL by retrieving relevant driving experiences from a memory database, ensuring scenario-specific contextual alignment. MSIER Luo et al. (2024a) further enhances example selection through a Multimodal Supervised In-Context Examples Retrieval framework, leveraging a foundation MLLM scorer to evaluate both textual and visual relevance. Meanwhile, Raven Rao et al. (2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
4.4.2 Reasoning
@@
Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems. RAGAR Khaliq et al. (2024) introduces Chain of RAG and Tree of RAG to iteratively refine fact-checking queries and explore branching reasoning paths for more robust evidence generation.
VisDoM Suri et al. (2024) integrates CoT with evidence curation to ensure logical and contextual accuracy, while SAM-RAG Zhai (2024) employs reasoning chains alongside multi-stage answer verification to enhance the relevance, utility, and support of generated responses. Meanwhile, LDRE Yang et al. (2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
4.4.3 Instruction Tuning
Several works have fine-tuned or instruct-tuned generation components for specific applications. RA-BLIP Ding et al. (2024b) leverages the Q-Former architecture from InstructBLIP Dai et al. (2023) to extract visual features based on question instructions, while RAGPT Lang et al. (2025) employs a context-aware prompter to generate dynamic prompts from relevant instances. mR2AG Zhang et al. (2024f) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks.
RagVL Chen et al. (2024d) employs instruction tuning to enhance the ranking capability of MLLMs, serving them as a re-ranker for filtering the top-k retrieved images. Jang et al. (2024) focus on distinguishing image differences to generate descriptive textual responses. MMed-RAG Xia et al. (2024a) applies preference fine-tuning to help models balance retrieved knowledge with internal reasoning.
To improve generation quality, MegaPairs Zhou et al. (2024a) and Surf Sun et al. (2024a) construct multimodal instruction-tuning datasets from prior LLM errors, while Rule Xia et al. (2024b) refines Med-LVLM through direct preference optimization to mitigate overreliance on retrieved contexts.
4.4.4 Source Attribution and Evidence Transparency
@@
Ensuring source attribution in multimodal RAG systems is a key focus of recent research. MuRAR Zhu et al. (2025) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring informativeness. VISA Ma et al. (2024b) uses large vision-language models to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots. Similarly, OMG-QA Nan et al. (2024) ensures transparency by prompting the LLM to explicitly cite evidence in generated responses.
@@ 4.5 Training Strategies
Training multimodal RAG models involves a multi-stage process to effectively handle cross-modal interactions Chen et al. (2022a). Pretraining establishes the foundation using large paired datasets to learn cross-modal relationships while fine-tuning adapts models to specific tasks by leveraging cross-modal attention Ye et al. (2019). For instance, REVEAL Hu et al. (2023) integrates multiple training objectives. Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where the model predicts text continuations from a prefix and an image. Supporting losses include Contrastive Loss ($L_{\text{contra}}$) for aligning queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to improve embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to align query and knowledge embeddings. During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning. Details of formulas for widely used RAG loss functions can be found in Appendix (§D).
4.5.1 Alignment
@@
The eCLIP loss Kumar and Marttinen (2024) extends contrastive training by integrating expert-annotated data and an auxiliary Mean Squared Error (MSE) loss to refine embedding quality. Mixup strategies further augment training by generating synthetic positive pairs, improving generalization in contrastive learning Kumar and Marttinen (2024). Dense2Sparse Nguyen et al. (2024) incorporates two unidirectional losses: an image-to-caption loss $\ell(I\to C)$ and a caption-to-image loss $\ell(C\to I)$. It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations.
4.5.2 Generation
A key aspect of multimodal RAG is the generation ability. Autoregressive language models are typically trained using Cross-Entropy Loss Brown et al. (2020). For image generation, widely used approaches include Generative Adversarial Networks (GANs) Goodfellow et al. (2014) and Diffusion Models Ho et al. (2020). GANs employ various loss functions, such as Binary Cross-Entropy Loss, Minimax Loss, Wasserstein Loss (WGAN), and Hinge Loss. Diffusion Models utilize Mean Squared Error (MSE) Loss for noise prediction, a common approach in Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020).
4.5.3 Robustness and Noise Management
@@
Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024). Managing noisy retrieval inputs is critical for maintaining model performance. MORE Cui et al. (2024) injects irrelevant results during training to enhance focus on relevant inputs. AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment. RAGTrans Cheng et al. (2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information. RA-BLIP Ding et al. (2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning. This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters.
RagVL Chen et al. (2024d) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise. Finally, RA-CM3 Yasunaga et al. (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.